The populalation of a toun grows at the rate of 20 persent every five years.In how much Years will it doubble itself? the answer is 20 plz help me finding the answer....Thanku
Answer by stanbon(75887)

The population of a town grows at the rate of 20 persent every five years.In how many Years will it double itself?
---------------------------
Let initial population be "x".
Solve x(1.20)^t = 2x
------
1.2^t = 2
t = ln(2)/ln(1.2)
time in 5-year periods = 3.8018
------
Time in one year periods = 3.8018*5 = 19.01 years
======================================================
Cheers,
Stan H.
===========